OpenStudy (anonymous):

Prove that, for all non-zero a and b,\[(ab)^x = a^x * b^x\] Prove that, for all non-zero a and b,\[(ab)^x = a^x * b^x\] @Mathematics

OpenStudy (jamesj):

Depends on your definition of \( x^y \).

OpenStudy (anonymous):

true, which is why I need help with that :(

OpenStudy (jamesj):

It's a bit messy and truly depends on where you are, so I'm quite serious as to what is your defintion of x^y

OpenStudy (anonymous):

I don't know the best definition... if I define it as a product using that big pi notation, I can't get all reals. If I use the exp() and ln() functions I only get the positive reals. I want all the reals to work.

OpenStudy (anonymous):

I'm sure there's a straightforward, albeit tedious, way to fit all the non-zero reals into the a and b, and all reals into the x

OpenStudy (asnaseer):

isn't if sufficient proof to just state:\[(ab)^x=(ab)*(ab)*(ab)*...\]so you get 'x' lots of multiplications and then just use the commutative and associative properties to re-arrange this as:\[=(a*a*a*...)*(b*b*b*...)\]and hence:\[=a^x*b^x\]

OpenStudy (asnaseer):

if == it

OpenStudy (anonymous):

what is '1/4' lots of multiplications, or '-1' lots if we do it that way? :(

OpenStudy (asnaseer):

hmmm... I see professor that you are playing with my mind :-)

OpenStudy (anonymous):

I will keep this question to annoy the math faculty.

OpenStudy (anonymous):

along with the torus cutting problem and some other ones

OpenStudy (jamesj):

In the most general form, we define exponent in the complex numbers and then restrict back to the real: Let z be a complex number and Arg(z) be defined as its principle argument, \(Arg(z) \in [0,2\pi) \) Then if ln is the usual real valued function for ln, we define the Ln(z) to be \[ Ln(z) = ln |z| + Arg(z) \] and another 'function', although it is actually not a function, \[ ln(z) = ln |z| + Arg(z) + 2k\pi, \ \ k \in Z \] Now, with that definition, we can make sense of \( w^z \) for any complex number \( z \) and \( w \neq 0 \): \[ w^z = exp(z . ln(w)) \] Notice by construction w^z has nearly always infinitely many values. We say the principle value is \[ exp(z . Ln(w)) \] Suppose now z and w are real numbers and w is not zero. Then w^z is the principle value of w^z considered as complex numbers provided that number is real. If it is not, then w^z is not defined. Finally, there is the issue of w = 0. Usually (but not always!) it is easiest to say that 0^z is zero provided Re(z) > 0 and undefined otherwise.

OpenStudy (jamesj):

Now with that definition, you can show your result if you're careful.

OpenStudy (anonymous):

I'll print that answer so once that Spivak book teaches me what a complex number is, my mind will be able to tackle this :-D

OpenStudy (asnaseer):

aha, could we use logs, let:\[n=(ab)^x\]\[\ln n=\ln (ab)^x=x\ln(ab)=x(\ln(a)+\ln(b))=x\ln(a)+x\ln(b)=\ln a^x+\ln b^x\]\[=\ln (a^x*b^x)\]therefore:\[n=a^x*b^x\]so:\[(ab)^x=a^x*b^x\]

OpenStudy (anonymous):

if we use logs, then we have problems when either a or b (and perhaps both at once) are negative

OpenStudy (jamesj):

When you do you'll see that exponents can become multi-valued very often and it leads to a radical change in the way we think of the domain of complex-valued functions, a topic called "Riemann surfaces".

OpenStudy (anonymous):

I will ask my calculus II professor what a Riemann surface is :-D

OpenStudy (asnaseer):

hmmm - this maths beast is a lot more subtle than I had envisaged :-( I guess you guys are working at a much higher plateau than I am at!

OpenStudy (jamesj):

lol. He's either going to love you or hate you.

OpenStudy (anonymous):

Alright; I can sleep now that I know that we can change (ab)^x into a^x * b^x